OpenStudy (anonymous):

solve this inequality and write the answer in interval notation: (4)/(x-3)-(1)/(2x)≥0

OpenStudy (anonymous):

How would you solve it if it were an equality?

OpenStudy (anonymous):

im not sure.

OpenStudy (anonymous):

(-infinity,-4)U(3/2,infinity)

OpenStudy (anonymous):

do you think you could show me how you got that answer?

OpenStudy (anonymous):

If you had: \[\frac{4}{x-3} - \frac{1}{2x} = 0\] What would your first step be.

OpenStudy (anonymous):

find a common denominator

OpenStudy (anonymous):

That's one way to start. Lets go with that since it's your intuition

OpenStudy (anonymous):

So what would that look like?

OpenStudy (anonymous):

i just want to know how you get the answer (-infinity,-4)U(3/2,infinity) ?

OpenStudy (anonymous):

This is how you get the answer

OpenStudy (anonymous):

Polpak is explaning how you get the answer. Listen, she/he knows what they are talking about. :)

OpenStudy (anonymous):

yes, how do i get the answer?

OpenStudy (anonymous):

The overall strategy is to solve the relation the same way you would solve an equality. And just remember if you multiply (or divide) by a negative you change the direction of the relation.

OpenStudy (anonymous):

ok but i still dont see how you get -4 and (3/2)

OpenStudy (anonymous):

Did you solve the relation for x?

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

-15

OpenStudy (anonymous):

That doesn't seem quite right. Care to show your work for that?

OpenStudy (anonymous):

can you just please show me how to solve it.